OpenStudy (anonymous):

MEDAL AND FAN AND GREAT TESTIMONIAL GIVEN :D PLEASE HELP :) starting with the second day of school, Jose and his friends traded stamps during recess and lunch. Jose got 4 new stamps that day. On the third day, he got 8 new stamps. On the fourth day, he got 12 new stamps. On the next day, he got 16 new stamps. If the pattern continues, how many new stamps will Jose get on the tenth day? How many new stamps will he have gotten in all?

OpenStudy (anonymous):

@satellite73 @mathslover @jim_thompson5910

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

any help on any of the attachments will be appreciated :) Thnx

OpenStudy (anonymous):

@satellite73 @saifoo.khan

jimthompson5910 (jim_thompson5910):

First term: 4 Common Difference: 8 - 4 = 4, 12 - 8 = 4, etc So a1 = 4 and d = 4 So you have an arithmetic sequence: a(n) = a1 + d(n-1) a(n) = 4 + 4(n-1) a(n) = 4 + 4n-4 a(n) = 4n

OpenStudy (ghazi):

tn= a + (n-1)d d=4, a= 12 and n=10

jimthompson5910 (jim_thompson5910):

the 10th term is a(n) = 4n a(10) = 4(10) a(10) = 40

jimthompson5910 (jim_thompson5910):

to find the sum of the first 10 terms, you would use this formula Sn = n*(a1 + an)/2

OpenStudy (anonymous):

wouldn't the 10th term equal 36?

jimthompson5910 (jim_thompson5910):

where n = 10, a1 = 4, an = 40 (tenth term)

OpenStudy (anonymous):

but it says 4 is in day 2? not day 1?

jimthompson5910 (jim_thompson5910):

oh true, so you would use n = 9 because n = 1 refers to day 2

OpenStudy (anonymous):

so it would be 36, or not?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

and the sum, would it be 180 in total?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

oh okaii thanx. can you help me with #2? maybe?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Ok I'll help with one more what did you get for # 2?

OpenStudy (anonymous):

oh one more, in that case i need more help in #4

jimthompson5910 (jim_thompson5910):

ok so do # 4 instead of # 2?

OpenStudy (anonymous):

yes please :)

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

sure

jimthompson5910 (jim_thompson5910):

ok so they switch when the total score is 5 or a multiple of it so if one side has 3 points and the other side has 2 points, then the total score is 3+2 = 5

jimthompson5910 (jim_thompson5910):

which means that the teams would change sides at this point

jimthompson5910 (jim_thompson5910):

if the score on one side is 6 and the other side is 4, then 6+4 = 10 the teams would flip sides at this point since 10 is a multiple of 5

OpenStudy (anonymous):

oh so the multiple would be if they add to 5 ?

jimthompson5910 (jim_thompson5910):

they switch yes, if the sum is 5, 10, 15, etc

OpenStudy (anonymous):

can you continue explaining the rest of this problem?

jimthompson5910 (jim_thompson5910):

they just want you to make up scores that add up to 5, 10, 15, etc

jimthompson5910 (jim_thompson5910):

3+2 = 5 6+14 = 10 are two sets of scores

jimthompson5910 (jim_thompson5910):

you could have 1, 4 ---> 1+4 =5 or 0 and 5 ---> 0+5 = 5 or any other combo that adds to 5

jimthompson5910 (jim_thompson5910):

then you just make up any final score that's reasonable (within the given limits of this problem)

OpenStudy (anonymous):

oh so the final scores have to add up to five?

jimthompson5910 (jim_thompson5910):

the final scores could be 15 to 12 since the first team reached 15

jimthompson5910 (jim_thompson5910):

another final score could be 27 to 25 since the first team scored 2 more points than the second

OpenStudy (anonymous):

but i thought it could not pass 15 since when they reach 15 they change?

jimthompson5910 (jim_thompson5910):

if one team reaches 15 then that teams wins if they have more than one point that the other team

jimthompson5910 (jim_thompson5910):

so 15 to 12 would have team A winning but 15 to 14 would mean the game isn't over yet

OpenStudy (anonymous):

yeah but you said "another final score could be 27 to 25 since the first team scored 2 more points than the second"?

jimthompson5910 (jim_thompson5910):

yes say the score is 26 to 25, the game isn't over but say team A scores one point to get 27 to 25, the game would be over with team A the winner

OpenStudy (anonymous):

oh but if they reach 15 they win, so how did they reach 27?

jimthompson5910 (jim_thompson5910):

if it's 15 to 14, the game isn't over

jimthompson5910 (jim_thompson5910):

15 to 15, game isn't over 15 to 16, game isn't over 16 to 16, game isn't over ... ... 25 to 25, game isn't over 26 to 25, game isn't over 27 to 25, team A wins

OpenStudy (anonymous):

oh okaii ... wait?

jimthompson5910 (jim_thompson5910):

what

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

but why would one team have 27 to 25 ? that is past the range of 15 for a team to win?

jimthompson5910 (jim_thompson5910):

you can only win if you get to 15 and your team has more than 1 point than the other

jimthompson5910 (jim_thompson5910):

15 to 14 would not end the game but 15 to 13 would (team A wins)

OpenStudy (anonymous):

yeah but u exceeded 15 when u got 27?

jimthompson5910 (jim_thompson5910):

yes i know but you need to be at least 2 points above the losing team

jimthompson5910 (jim_thompson5910):

technically, the game could go on forever if you keep tying

OpenStudy (anonymous):

so you win if one team get like 15 to 13 right? or like 12 to 10? or even like 22 to 25? am i right?

jimthompson5910 (jim_thompson5910):

you have to be at least 2 pts above and be at 15 or greater

jimthompson5910 (jim_thompson5910):

12 to 10 the games not over yet

OpenStudy (anonymous):

wait why?

jimthompson5910 (jim_thompson5910):

because neither team has gotten to 15 yet

jimthompson5910 (jim_thompson5910):

22 to 25 would have team B win, but that's not a possible score

jimthompson5910 (jim_thompson5910):

22 to 24 would work though

OpenStudy (anonymous):

why would it not work?

jimthompson5910 (jim_thompson5910):

because 22 to 24 would end the game

jimthompson5910 (jim_thompson5910):

22 to 25 isn't possible to reach

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

is it cause its +3 more?

jimthompson5910 (jim_thompson5910):

yeah but once the score is 22 to 24, the game is over

OpenStudy (anonymous):

oh okaii got you so can you help me with #2 please? that is all i need?

jimthompson5910 (jim_thompson5910):

no sry, i have to get going

OpenStudy (anonymous):

oh, okaii r u going to help others now?

OpenStudy (anonymous):

thanx though really appreciate it :)

OpenStudy (anonymous):

@satellite73 @ghazi @mathslover @abb0t

OpenStudy (anonymous):

@satellite73 @sammie12